OpenStudy (anonymous):

If a ball is thrown vertically upward from the roof of 64 foot building with a velocity of 64 ft/sec, then what is the maximum height the ball reaches? What is the velocity of the ball when it hits the ground?

OpenStudy (anonymous):

Set the derivative of the equation equal to zero to find the maximum height ds/dt = 16-32t= 0 t = 16/32 = .5 seconds substitute that in the equation s=64+16(0.5)-16(0.5)^2 to find part b you set s(t)=0 then solve the quadratic equation to get two values of t ignore the negative one, then u know that V(t) is derivative of s(t) so u find the derivative and substitute the time u get, to find the velocity first set s(t) =0 so u have 64+16t−16t^2=0 now solve this quadratic equations u will get two solutions for t u should get 7.5 and -0.5

OpenStudy (anonymous):

I plugged in 7.5 for maximum height and it doesnt work

OpenStudy (anonymous):

@im.celibate

OpenStudy (anonymous):

Oh. That is weird because I got the answer right just like 2 hours ago..

OpenStudy (anonymous):

i plugged it online it does not wrk

OpenStudy (anonymous):

oh, that sucks

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

Use the formula: v^2 = u^2 + 2as Where the final velocity is 0 (maximum height: think of it...) initial velocity = 64, acceleration = 32.2 and s = displacement (your maximum height)

OpenStudy (anonymous):

So: 0^2 = 64^2 + 2 x 32.2 x s Now solve for s